OpenStudy (anonymous):

Evaluate the integral as a series. \[\int xcos(x^3)dx\] \[cosx=\sum_{n=0}^{\infty}(-1)^n\frac{x^{2n}}{(2n)!}\] sooooo..... \[=\int\sum_{n=0}^{\infty}(-1)^n\frac{x^{6n+1}}{(2n)!}\] Yes?

OpenStudy (rogue):

yup, now just integrate.

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}\frac{(-1)^n}{(2n)!}\frac{x^{6n+2}}{6n+2}\]

OpenStudy (anonymous):

Awww thank you!